### v1

@@ 1. Introduction
Let $\rho:\Gamma_{g}\to\operatorname{SL}(2,\mathbb{C})$ is nearly Fuchsian, i.e. suppose that it admits an equivariant immersion $h:\widetilde{S_{g}}\to\mathbb{H}^{3}$ with principal curvature in $(-1,1)$. Epstein showed that the locally symmetric space $\mathbb{H}^{3}/\rho(\Gamma_{g})$ admits a fibration described by the fitting immersion $\mathcal{G}h$ that associates to $x\in\widetilde{S_{g}}$ the geodesic orthogonal to $h(\widetilde{S_{g}})$ at $h(x)$ [Eps86]. The map $\mathcal{G}h$ is sometimes referred to as the Gauss map of $h$.
@@
In order to show this result we introduce the notion of a fitting flow. Let $\mathcal{E}$ be the tautological rank $2$ vector bundle over $\text{Gr}_{2}(\mathcal{Q})$ and $\pi:\mathcal{E}\to\mathcal{Q}$ be the tautological projection. The pullback $u^{*}\mathbb{S}\mathcal{E}$ of the circle bundle $\mathbb{S}\mathcal{E}$ for a map $u:S\to\text{Gr}^{\text{mix}}_{2}(\mathcal{Q})$ is the space of pairs $(x,H)$ where $x\in S$ and $H$ is a co-oriented projective hyperplane in $\mathbb{P}(S^{2}V)$ containing the codimension $2$ projective subspace associated to $u(x)$. Note that a co-oriented projective hyperplane intersecting $\mathbb{P}(S^{2}V^{>0})$ defines a half space in the convex set $\mathbb{P}(S^{2}V^{\geq 0})$
@@
In order to construct continuous maps that admits a fitting flow for maximal representations, we show how to associate to a pair of transverse Lagrangians a quadric on $\mathbb{R}^{2n}$. We discuss an other use for these quadrics in Section LABEL:subsec:polygons_of_hyperboloids to decompose the projective structures associated to maximal representations into smaller pieces which are polygons of quadric hypersurfaces. These quadrics were used by Burelle-Treib to study Schotty subgroups [BT17].
@@ Acknowledgments
I would like to thank Andrea Seppi for the helpful discussions during my stay in Grenoble that helped me a lot on this project. I would like to thank also beatrice Pozzetti for discussing this topic and giving feedback. The author was funded through the DFG Emmy Noether project 427903332 of B. Pozzetti and the RTG 2229 “Asymptotic Invariants and Limits of Groups and Spaces”.
@@
The sum of the $(\epsilon_{i})$ is an invariant of the triple of flags that is called the Maslov index $M(\ell_{1},\ell_{2},\ell_{3})$. The group $\operatorname{Sp}(2n,\mathbb{R})$ acts transitively on the space of triples of transverse Lagrangians with a given Maslov index. We say that $(\ell_{1},\ell_{2},\ell_{3})$ is maximal if the Maslov index of the triple in equal to $n$.
The Lie group $\operatorname{Sp}(2n,\mathbb{R})$ is of Hermitian type and tube type. Hence it admits a special class in its continuous cohomology group $[\tau]\in H^{2}_{c}(\operatorname{Sp}(2n,\mathbb{R}),\mathbb{Z})$. Let $S_{g}$ be a closed oriented surface of genus $g\geq 2$. The fundamental class of $S_{g}$ defines a cohomology class $[S_{g}]\in H^{2}(\pi_{1}(S_{g}),\mathbb{Z})\simeq\mathbb{Z}$. Given a representation $\rho:\pi_{1}(S_{g})\to\operatorname{Sp}(2n,\mathbb{R})$ one can consider the pullback of this class $\rho^{*}[\tau]=T(\rho)[S_{g}]$. the integer $T(\rho)$ is called the Toledo number of $\rho$.
@@
Such a representation is called maximal if its Toledo number is equal to $(2g-2)n$.
@@
Fix a word metric ||$\cdot$|| ⋅ | on $\Gamma$ and a scalar product on $\mathbb{R}^{2n}$ allowing us to define the singular values $(\sigma_{1}(g)\geq\sigma_{2}(g)\geq\cdots\geq\sigma_{2n}(g)$ of $g\in\operatorname{SL}(2n,\mathbb{R})$ as the eigenvalues of $\sqrt{g^{t}g}$. The following definition is independent of these choices.
@@
A representation $\rho:\pi_{1}(S_{g})\to\operatorname{PSp}(2n,\mathbb{R})$ is maximal if and only if it is $\{n\}$-Anosov and for one and hence any positively oriented triple $(x,y,z)\in\partial\pi_{1}(S)$ the triple $(\xi^{n}_{\rho}(x),\xi^{n}_{\rho}(y),\xi^{n}_{\rho}(z))$ is a maximal triple of Lagragians.
@@
In Theorem 2.4 and hence also in Theorem 2.5, one can relax the assumption that $\rho$ is $\{n\}$-Anosov. It is sufficient to assume that $\rho$ admits an equivariant, continuous and transverse map $\xi^{n}_{\rho}:\partial\Gamma\to\mathcal{L}_{n}$ satisfying the additional property.
@@
Let $\rho:\pi_{1}(S_{g})\to\operatorname{PSp}(2n,\mathbb{R})$ be $\{n\}$-Anosov. Let $(x,y,z)$ be a positively oriented triple in $\partial\pi_{1}(S_{g})$. Up to changing the symplectic basis, we can assume that for some $(\epsilon_{i})\in\{-1,1\}$:
@@
These loops are homotopic to $\tau$ or its inverse depending on the sign of $\epsilon_{i}$. The homotopy type of $\tau_{0}$ is hence equal to $(\epsilon_{1}+\epsilon_{2}+\cdots+\epsilon_{n})[\tau]$. Moreover the set of Lagrangians tranverse to a given Lagrangian is contractible, so one can homotope $\tau_{0}$ on the intervals $[0,\frac{\pi}{2}]$, $[\frac{\pi}{2},\pi]$ and $[\pi,2\pi]$ to coincide with $\xi^{n}_{\rho}$. Hence the free homotopy class of $\xi^{n}_{\rho}$ is equal to the one of $\tau_{0}$, which is equal to $M(\xi^{n}_{\rho}(x),\xi^{n}_{\rho}(y),\xi^{n}_{\rho}(z))[\tau]$. We therefore deduce that $\rho$ is maximal if and only if it is $\{n\}$-Anosov and $[\xi^{n}_{\rho}]=n[\tau]$. ∎
@@ 3.1. Pencils of quadrics
Let $V$ be a finite even-dimensional vector space. Let $S^{2}V$ be the space of symmetric bilinear tensors on $V$, which we interpret as maps $V^{*}\to V$. The dual space $\mathcal{Q}=S^{2}V^{*}$ is the space of symmetric bilinear forms on $V$, or the space of quadrics on $V$, that we interpret as maps $V\to V^{*}$.
We denote by $S^{2}V^{\geq 0}$ and $S^{2}V^{>0}$ respectively the space of semi-positive and positive symmetric tensors, i.e. elements $p\in S^{2}V$ such that $p^{-1}$ is respectively a semi-positive and positive bilinear form. The Lie group $\operatorname{SL}(V)$ acts on $S^{2}V$, and preserves the properly convex set $\mathbb{P}(S^{2}V^{\geq 0})$. The convex domain $\mathbb{P}(S^{2}V^{>0})$ is a projective model for the symmetric space associated to $\operatorname{SL}(V)$.
@@
To an element $P\in\operatorname{Gr}_{d}(\mathcal{Q})$ one can associate its annihilator codimension $d$ subspace $P^{\circ}\subset S^{2}V$. This dual space can be described as the space of symmetric tensors $p$ on which one has $q(p)=\operatorname{Tr}(q\circ p)=0$ for all $q\in P$. Note that the projectivization $\mathbb{P}(P^{\circ})$ also has codimension $d$ in $\mathbb{P}(S^{2}V)$.
This subspace does not necessarily intersect the convex $\mathbb{P}(S^{2}V^{>0})$.
@@
Equivalently $P$ is mixed if and only if it does not contain any semi-positive quadric $0\neq q\in P$. Indeed the dual of the cone of positive elements $S^{2}V^{>0}$ is the cone of semi-positive bilinear forms in $\mathcal{Q}$.
@@
To a pencil $P\in\operatorname{Gr}_{d}(\mathcal{Q})$ we associate the codimension $d$ subspace $\mathbb{P}(P^{\circ}\cap S^{2}V^{\geq 0})$ in the convex set $\mathbb{P}(S^{2}V^{\geq 0})$.
We say that two elements $P_{1},P_{2}\in\operatorname{Gr}_{d}(\mathcal{Q})$ form a fitting pair if the associated subspaces $\mathbb{P}(P_{1}^{\circ}\cap S^{2}V^{\geq 0})$ and $\mathbb{P}(P_{2}^{\circ}\cap S^{2}V^{\geq 0})$ are disjoint.
The structure of the convex set $\mathbb{P}(S^{2}V^{\geq 0})$ is involved, but the set of its extremal points $S^{2}\mathbb{P}(V)$ is the projectivization of the set of rank one tensors $S^{2}\mathbb{P}(V)$, which is in one-to-one correspondence with $\mathbb{P}(V)$. We show that the condition of being a fitting pair can be checked by looking only at $\mathbb{P}(V)$.
@@
Let $u:M\to\operatorname{Gr}^{{\operatorname{mix}}}_{d}(\mathcal{Q})$ be continuous, the following are equivalent:
@@
In order to prove this, note that the convex set $\mathbb{P}(S^{2}V^{\geq 0})$ does not have any segment as a facet, i.e. the intersection of the convex with a supporting hyperplane. Hence it satisfies the following lemma.
@@
For all linear hyperplane $H$ in $S^{2}V$ the extremal points of $\mathbb{P}(H\cap S^{2}V^{\geq 0})$ are also extremal points of $\mathbb{P}(S^{2}V^{\geq 0})$. In particular:
@@
Let $f$ be a facet of $\mathbb{P}(S^{2}V^{\geq 0})$. It is the intersection of this convex with a projective hyperplane corresponding to $[q]\in\mathbb{P}(S^{2}V^{*})=\mathbb{P}(\mathcal{Q})$. The fact that this is a supporting hyperplanes implies that $q\in\mathcal{Q}$ is a semi-positive element. Let $W\subset V$ be the vector subspace of isotropic vectors for $q$. The corresponding facet is equal to $\mathbb{P}(S^{2}W^{\geq 0})$.
Hence facets of $\mathbb{P}(S^{2}V^{\geq 0})$ are of the form $\mathbb{P}(S^{2}W^{\geq 0})$ for $W\subset V$ a linear subspace. This has dimension 00 or at least $2$, and therefore no facet is a segment.
Suppose that for some general projective hyperplane, some extremal point $p$ of $\mathbb{P}(H\cap S^{2}V^{\geq 0})$ is not an extremal point of $\mathbb{P}(S^{2}V^{\geq 0})$, then it belongs to the interior of a facet of $\mathbb{P}(S^{2}V^{\geq 0})$, that has dimension at least $2$ as stated previously. The intersection of this face with $H$ contains therefore a segment, so $p$ is not an extremal point of $H\cap\mathbb{P}(S^{2}V^{\geq 0})$. ∎
@@
The space $\operatorname{Gr}_{d}(\mathcal{Q})$ inherits the structure of a smooth manifold. A chart around a point $P\in\operatorname{Gr}_{d}(\mathcal{Q})$ can be constructed given a subspace $Q$ such that $P\oplus Q=V$. We denote by $U_{Q}\subset\operatorname{Gr}_{d}(\mathcal{Q})$ the open set of elements transverse to $Q$. Every element of $U_{Q}$ can be written uniquely as the graph $\{x+u(x)|x\in P\}$ for some linear map $u:P\to Q$. Hence $U_{Q}$ can be identified with the vector space $\operatorname{Hom}(P,Q)$.
@@
Let $\mathrm{v}\in T_{P}\operatorname{Gr}_{d}(\mathcal{Q})$, that we see as an element of $\operatorname{Hom}(P,\mathcal{Q}/P)$. One can interpret $\operatorname{Ker}(\mathrm{v})$ as the set of vectors in $P$ which remains at first order in the dimension $d$ linear subspace when the subspace moves in the direction $\mathrm{v}$. More precisely the lines $\ell$ in the kernel are exactly the one such that for any Riemannian metric on $\mathbb{P}(V)$ and every curve $\gamma$ in $\operatorname{Gr}_{P}(V)$ with $\gamma(0)=P$ and $\gamma^{\prime}(0)=\mathrm{v}$ one has $d\left(\mathbb{P}\left(\gamma(t)\right),\ell\right)=o(t)$ at $t=0$. Because of the identification $T_{P}\operatorname{Gr}_{d}(\mathcal{Q})\simeq T_{P^{\circ}}\operatorname{Gr}_{N-d}(\mathcal{Q}^{*})$ where $N=\dim(\mathcal{Q})$, to such an element $\mathrm{v}$ corresponds an element:
@@
We check that these two statements are indeed equivalent. We first show that $\operatorname{Ker}(\mathrm{v}^{\circ})=\operatorname{Im}(\mathrm{v})^{\circ}$, where we use the natural identification $(\mathcal{Q}/P)^{*}\simeq P^{\circ}$.
We prove this by writing an equation that relates $\mathrm{v}$ and $\mathrm{v}^{\circ}$. By definition, $\operatorname{Tr}(qp)=0$ for $q\in P$ and $p\in P^{\circ}$. Let us choose some representatives $\overline{\mathrm{v}(p)}\in S^{2}V$ and $\overline{\mathrm{v}^{\circ}(q)}\in\mathcal{Q}$ for $\mathrm{v}(p)\in S^{2}V/P^{\circ}$ and $\mathrm{v}^{\circ}(q)\in\mathcal{Q}/P$. Let us fix $q\in P$ and $p\in P^{\circ}$. If $(P_{t})$ is a smooth curve with $P_{0}=P$ and with derivative $\mathrm{v}$ at $t=0$, $p+t\overline{\mathrm{v}^{\circ}(p)}+o(t^{2})\in P_{t}$ and $q+t\overline{\mathrm{v}(q)}+o(t^{2})\in P^{\circ}_{t}$. Hence we get:
$\operatorname{Tr}\left(q\overline{\mathrm{v}^{\circ}(p)}+\overline{\mathrm{v}(q)}p\right)=0.$
An element $p\in P^{\circ}$ satisfies $p\in\operatorname{Ker}(\mathrm{v}^{\circ})$ if and only if $\operatorname{Tr}(\overline{\mathrm{v}(q)}p)=0$ for all $q\in\mathcal{Q}$, hence if and only if the corresponding linear form on $Q/P$ belongs to $\operatorname{Im}(\mathrm{v})^{\circ}$.
Now we prove the equivalence of the two definitions. If there exist $[q]\in\operatorname{Im}(\mathrm{v})$ with $q$ positive, then for any $p\in\operatorname{Ker}(\mathrm{v}^{\circ})$, $\operatorname{Tr}(pq)=0$. Hence $p$ is not a positive tensor. Therefore $\operatorname{Ker}(\mathrm{v}^{\circ})\subset P^{\circ}$ intersects trivially $S^{2}V^{\geq 0}$.
Conversely if $\operatorname{Ker}(\mathrm{v}^{\circ})$ intersects trivially $S^{2}V^{\geq 0}$, there exist $q\in\mathcal{Q}$ that do not vanish on $S^{2}V^{\geq 0}$, i.e. $q$ is positive. The class $[q]\in\mathcal{Q}/P$ belongs to $\operatorname{Im}(\mathrm{v})$.

### v2

@@ 1. Introduction
[…]
@@
As a corollary we prove that there exist quasi Fuchsian representations that are not nearly Fuchsian.
@@
In order to show this result we introduce the notion of a fitting flow. Let $\mathcal{E}$ be the tautological rank $2$ vector bundle over $\text{Gr}_{2}(\mathcal{Q})$, i.e. :
$\mathcal{E}=\{(P,q)\mid P\in\text{Gr}_{2}(\mathcal{Q}),\,q\in P\}.$
Let $\pi:\mathcal{E}\to\mathcal{Q}$ be the tautological projection, i.e. the projection onto the second factor. The pullback $u^{*}\mathbb{S}\mathcal{E}$ of the circle bundle $\mathbb{S}\mathcal{E}$ for a map $u:S\to\text{Gr}^{\text{mix}}_{2}(\mathcal{Q})$ is the space of pairs $(x,H)$ where $x\in S$ and $H$ is a co-oriented projective hyperplane in $\mathbb{P}(S^{2}V)$ containing the codimension $2$ projective subspace associated to $u(x)$. Note that a co-oriented projective hyperplane intersecting $\mathbb{P}(S^{2}V^{>0})$ defines a half space in the convex set $\mathbb{P}(S^{2}V^{\geq 0})$
@@
In order to construct continuous maps that admits a fitting flow for maximal representations, we show how to associate to a pair of transverse Lagrangians a quadric on $\mathbb{R}^{2n}$.
@@ Acknowledgments
I would like to thank Andrea Seppi for the helpful discussions during my stay in Grenoble that helped me a lot on this project. I would like to thank also Beatrice Pozzetti for discussing this topic and giving feedback. The author was funded through the DFG Emmy Noether project 427903332 of B. Pozzetti and the RTG 2229 “Asymptotic Invariants and Limits of Groups and Spaces”, and partially by the European Union via the ERC 101124349 "GENERATE".
Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.
@@
The sum of the $(\epsilon_{i})$ is an invariant of the triple of flags that is called the Maslov index $M(\ell_{1},\ell_{2},\ell_{3})$. These facts can be found in [LV13, Section 1.5.7]. The group $\operatorname{Sp}(2n,\mathbb{R})$ acts transitively on the space of triples of transverse Lagrangians with a given Maslov index. We say that $(\ell_{1},\ell_{2},\ell_{3})$ is maximal if the Maslov index of the triple in equal to $n$.
The Lie group $\operatorname{Sp}(2n,\mathbb{R})$ is of Hermitian type and tube type. Hence it admits a special class in its continuous cohomology group $\tau]\in H^{2}_{c}(\operatorname{Sp}(2n,\mathbb{R}),\mathbb{Z})$. Let $S_{g}$ be a closed oriented surface of genus $g\geq 2$. The fundamental class of $S_{g}$ defines a cohomology class $S_{g}]\in H^{2}(\pi_{1}(S_{g}),\mathbb{Z})\simeq\mathbb{Z}$. Given a representation $\rho:\pi_{1}(S_{g})\to\operatorname{Sp}(2n,\mathbb{R})$ one can consider the pullback of this class $\rho^{*}[\tau]=T(\rho)[S_{g}]$. the integer $T(\rho)$ is called the Toledo number of $\rho$.
@@
Such a representation is called maximal if its Toledo number is equal to (2g2)n(2g-2)n.
@@
Fix a word metric $|\cdot|$ on $\Gamma$ and a scalar product on $\mathbb{R}^{2n}$ allowing us to define the singular values $\sigma_{1}(g)\geq\sigma_{2}(g)\geq\cdots\geq\sigma_{2n}(g)$ of $g\in\operatorname{SL}(2n,\mathbb{R})$ as the eigenvalues of $\sqrt{g^{t}g}$. The following definition is independent of these choices.
@@
The fact that $\xi^{n}_{\rho}(x)$ is a Lagrangian is a consequence of the fact that an attracting fixed $n$-dimensional subspace for an element $g\in\operatorname{Sp}(2n,\mathbb{R})$ is necessarily Lagrangian, and every $x\in\partial\Gamma$ is a limit of attracting fixed points $\gamma^{+}$ of elements $\gamma\in\Gamma$.
@@
A representation $\rho:\pi_{1}(S_{g})\to\operatorname{Sp}(2n,\mathbb{R})$ is maximal if and only if it is $\{n\}$-Anosov and for one and hence any positively oriented triple $(x,y,z)\in\partial\pi_{1}(S)$ the triple $(\xi^{n}_{\rho}(x),\xi^{n}_{\rho}(y),\xi^{n}_{\rho}(z))$ is a maximal triple of Lagragians.
@@
In Theorem 2.4 and hence also in Theorem 2.5, one can relax the assumption that $\rho$ is $\{n\}$-Anosov. In [BIW03, Theorem 8] is is shown that it is sufficient to assume that $\rho$ admits an equivariant, continuous and transverse map $\xi^{n}_{\rho}:\partial\Gamma\to\mathcal{L}_{n}$ satisfying the additional property of sending positive triples to maximal triples.
@@
Let $\rho:\pi_{1}(S_{g})\to\operatorname{Sp}(2n,\mathbb{R})$ be $\{n\}$-Anosov. Let $(x,y,z)$ be a positively oriented triple in $\partial\pi_{1}(S_{g})$. Up to changing the symplectic basis, we can assume that for some $(\epsilon_{i})\in\{-1,1\}$:
@@
These loops are homotopic to $\tau$ or its inverse depending on the sign of $\epsilon_{i}$. The homotopy type of $\tau_{0}$ is hence equal to $(\epsilon_{1}+\epsilon_{2}+\cdots+\epsilon_{n})[\tau]$. Moreover the set of Lagrangians tranverse to a given Lagrangian is contractible, so one can homotope $\tau_{0}$ on the intervals $0,\frac{\pi}{2}]$, $\frac{\pi}{2},\pi]$ and $\pi,2\pi]$ to coincide with $\xi^{n}_{\rho}$. Hence the free homotopy class of $\xi^{n}_{\rho}$ is equal to the one of $\tau_{0}$, which is equal to $M(\xi^{n}_{\rho}(x),\xi^{n}_{\rho}(y),\xi^{n}_{\rho}(z))[\tau]$. We therefore deduce that $\rho$ is maximal if and only if it is $\{n\}$-Anosov and $\xi^{n}_{\rho}]=n[\tau]$. ∎
@@ 3.1. Pencils of quadrics
Let $V$ be a finite even-dimensional vector space with a fixed volume form. Let $S^{2}V$ be the space of symmetric bilinear tensors on $V$, which we interpret as maps $V^{*}\to V$. The dual space $\mathcal{Q}=S^{2}V^{*}$ is the space of symmetric bilinear forms on $V$, or the space of quadrics on $V$, that we interpret as maps $V\to V^{*}$.
We denote by $S^{2}V^{\geq 0}$ and $S^{2}V^{>0}$ respectively the space of semi-positive and positive symmetric tensors, i.e. elements $s\in S^{2}V$ that define respectively a semi-positive and positive symmetric bilinear form on $V^{*}$. The Lie group $\operatorname{SL}(V)$ acts on $S^{2}V$, and preserves the properly convex set $\mathbb{P}(S^{2}V^{\geq 0})$. The convex domain $\mathbb{P}(S^{2}V^{>0})$ is a projective model for the symmetric space associated to $\operatorname{SL}(V)$.
@@
To an element $P\in\operatorname{Gr}_{d}(\mathcal{Q})$ one can associate its annihilator, or dual, codimension $d$ subspace $P^{\circ}\subset S^{2}V$. This dual space can be described as the space of symmetric tensors $s\in S^{2}V$ on which one has $q(s)=\operatorname{Tr}(q\circ s)=0$ for all $q\in P$. Note that the projectivization $\mathbb{P}(P^{\circ})$ also has codimension $d$ in $\mathbb{P}(S^{2}V)$.
The subspace $\mathbb{P}(P^{\circ})$ does not necessarily intersect the convex $\mathbb{P}(S^{2}V^{>0})$.
@@
Proposition 3.2.
A pencil $P\in\operatorname{Gr}_{d}(\mathcal{Q})$ is mixed if and only if there is no semi-positive quadric $0\neq q\in P$.
Morally this proposition is a consequence of the fact that the dual of the cone of positive elements $S^{2}V^{>0}$ is the cone of semi-positive bilinear forms in $\mathcal{Q}$, in the sense that $S^{2}V^{>0}$ is the set of tensors $s$ such that $q(s)>0$ for all semi-positive $q\in\mathcal{Q}$.
Proof.
Fix $P\in\operatorname{Gr}_{d}(\mathcal{Q})$ and suppose that there is a positive element $s\in E^{\circ}$. We write $s$ as a finite sum of positive rank one elements of the form $v_{i}\otimes v_{i}$ with $v_{i}\in V$ non-zero for $i\in I$ in a finite set. Indeed $s$ defines a positive bilinear form on $V^{*}$ and hence is a sum of rank one symmetric positive bilinear forms. Let $q\in P$ be a semi-positive element. One has $q(s)=0$ since $s\in P^{\circ}$ but $q(v_{i}\otimes v_{i})=q(v_{i},v_{i})\geq 0$ since $q$ is positive and hence for all $i\in I$, $q(v_{i},v_{i})=0$.
Note that for all $v\in V$ non-zero we can chose a decomposition of $s$ such that $v$ is one of the $v_{i}$ , since $s-\epsilon v\otimes v$ is still positive for $\epsilon$ small enough.
We proved that $q(v,v)=0$ for all $v\in V$ so $q$ is necessarily equal to 0. Hence $P$ does not contain any non-zero semi-positive element.
Now suppose that $P^{\circ}$ is disjoint from the cone $S^{2}V^{>0}$, so there must be a hyperplane $H$ in $S^{2}V$ containing $P^{\circ}$ and disjoint from $S^{2}V^{>0}$. This hyperplane corresponds to $\langle q\rangle^{\circ}$ for some non-zero $q\in P$. This element has the property that $q(s)\neq 0$ for all $s\in S^{2}V^{>0}$. Up to exchanging $q$ by q-q one can assume that $q(s)>0$ on $S^{2}V^{>0}$, so $q(s)\geq 0$ on $S^{2}V^{\geq 0}$. In particular for all $v\in V$, one has $q(v\otimes v)=q(v,v)\geq 0$, so $q$ is semi-positive.
@@
To a pencil $P\in\operatorname{Gr}_{d}(\mathcal{Q})$ we associate the codimension $d$ projective subspace $\mathbb{P}(P^{\circ})$, and its intersection $\mathbb{P}(P^{\circ}\cap S^{2}V^{\geq 0})$ in the convex set $\mathbb{P}(S^{2}V^{\geq 0})$.
Definition 3.3.
We say that two elements $P_{1},P_{2}\in\operatorname{Gr}^{\text{mix}}_{d}(\mathcal{Q})$ form a fitting pair if the associated subspaces $\mathbb{P}(P_{1}^{\circ}\cap S^{2}V^{\geq 0})$ and $\mathbb{P}(P_{2}^{\circ}\cap S^{2}V^{\geq 0})$ are disjoint.
The structure of the convex set $\mathbb{P}(S^{2}V^{\geq 0})$ is involved, but the set of its extreme points $S^{2}\mathbb{P}(V)\subset\mathbb{P}(S^{2}V^{\geq 0})$ is the projectivization of the set of rank one tensors, which is in one-to-one correspondence with $\mathbb{P}(V)$. Here the rank of a symmetric tensor is also the rank of the corresponding map $V^{*}\to V$. The extreme points of a closed properly convex subset $C$ of projective space are the points $x\in C$ such that if $s\subset C$ is a projective segment containing it, $x$ is an endpoint of $s$. We show that the condition of being a fitting pair can be checked by looking only at $\mathbb{P}(V)$.
@@
Let $P_{1},P_{2}\in\operatorname{Gr}^{{\operatorname{mix}}}_{d}(\mathcal{Q})$ be two mixed pencils.
@@
In order to prove this statement, the key fact that we will prove and use the is that the convex set $\mathbb{P}(S^{2}V^{\geq 0})$ does not have any segment as a face, i.e. the intersection of this convex with any hyperplane intersecting trivially the interior $\mathbb{P}(S^{2}V^{>0})$ is never a non-trivial segment.
@@
For all linear hyperplane $H$ in $S^{2}V$ the extreme points of $\mathbb{P}(H\cap S^{2}V^{\geq 0})$ are also extreme points of $\mathbb{P}(S^{2}V^{\geq 0})$. In particular:
@@
Recall that the Krein–Milman theorem states that the convex Hull of the extreme points of a closed convex set $C$ in a finite dimensional vector space is equal to $C$.
@@
Let $f$ be a face of $\mathbb{P}(S^{2}V^{\geq 0})$, i.e. the intersection of $\mathbb{P}(S^{2}V^{\geq 0})$ with a projective hyperplane corresponding to $q]\in\mathbb{P}(S^{2}V^{*})=\mathbb{P}(\mathcal{Q})$ not intersecting $\mathbb{P}(S^{2}V^{>0})$ trivially. It is the intersection of this convex with a projective hyperplane corresponding to $q]\in\mathbb{P}(S^{2}V^{*})=\mathbb{P}(\mathcal{Q})$. The fact that this hyperplanes intersects $\mathbb{P}(S^{2}V^{>0})$ trivially implies that there is a lift $q\in\mathcal{Q}$ that is a semi-positive element, by Proposition 3.2. Let $W\subset V$ be the vector subspace of isotropic vectors for $q$. The corresponding face is equal to $\mathbb{P}(S^{2}W^{\geq 0})$.
Hence faces of $\mathbb{P}(S^{2}V^{\geq 0})$ are of the form $\mathbb{P}(S^{2}W^{\geq 0})$ for $W\subset V$ a linear subspace. This has dimension 0 or at least $2$, and therefore no face is a segment.
Suppose that for some general projective hyperplane, some extreme point $[s]$ of $\mathbb{P}(H\cap S^{2}V^{\geq 0})$ is not an extreme point of $\mathbb{P}(S^{2}V^{\geq 0})$, then it belongs to the interior of a face of $\mathbb{P}(S^{2}V^{\geq 0})$, that has dimension at least $2$ as stated previously. The intersection of this face with $H$ contains therefore an open segment containing $s$, so $s$ is not an extreme point of $H\cap\mathbb{P}(S^{2}V^{\geq 0})$. ∎
@@
The space $\operatorname{Gr}_{d}(\mathcal{Q})$ inherits the structure of a smooth manifold. A chart around a point $P\in\operatorname{Gr}_{d}(\mathcal{Q})$ can be constructed given a subspace $Q$ such that $P\oplus Q=\mathcal{Q}$. We denote by $U_{Q}\subset\operatorname{Gr}_{d}(\mathcal{Q})$ the open set of elements transverse to $Q$. Every element of $U_{Q}$ can be written uniquely as the graph $\{x+u(x)|x\in P\}$ for some linear map $u:P\to Q$. Hence $U_{Q}$ can be identified with the vector space $\operatorname{Hom}(P,Q)$.
@@
Because of the identification $T_{P}\operatorname{Gr}_{d}(\mathcal{Q})\simeq T_{P^{\circ}}\operatorname{Gr}_{N-d}(\mathcal{Q}^{*})$ where $N=\dim(\mathcal{Q})$, to such an element $\mathrm{v}$ corresponds an element:
@@
Let $\mathrm{v}^{\circ}\in T_{P}\operatorname{Gr}_{d}(\mathcal{Q}^{*})$, that we see as an element of $\operatorname{Hom}(P^{\circ},\mathcal{Q}/P^{\circ})$. Let $(P^{\circ}_{t})_{t\geq 0}$ a curve of elements of $\operatorname{Gr}_{N-d}(\mathcal{Q}^{*})$ with $P^{\circ}_{0}=P^{\circ}$ and derivative $\mathrm{v}^{\circ}$ at $t=0$. One can interpret $\operatorname{Ker}(\mathrm{v}^{\circ})$ as the set of vectors in $P^{\circ}$ which remain in $P^{\circ}_{t}$ at first order around $t=0$. More precisely the lines $[s]$ in $\operatorname{Ker}(\mathrm{v}^{\circ})$ are exactly the ones such that for any Riemannian metric on $\mathbb{P}(V)$ and any such curve $(P^{\circ}_{t})_{t>0}$ one has $d\left(\mathbb{P}\left(P^{\circ}_{t}\right),[s]\right)=o(t)$ close to $t=0$.
@@
We now check that these two statements in the definition are indeed equivalent.
@@
We first show that $\operatorname{Ker}(\mathrm{v}^{\circ})=\operatorname{Im}(\mathrm{v})^{\circ}$, where we use the natural identification $(\mathcal{Q}/P)^{*}\simeq P^{\circ}$.
We prove this by writing an equation that relates $\mathrm{v}$ and $\mathrm{v}^{\circ}$. By definition, $\operatorname{Tr}(qs)=0$ for $q\in P$ and $s\in P^{\circ}$. Let us fix $q\in P$ and $s\in P^{\circ}$ and choose some representatives $\overline{\mathrm{v}(s)}\in S^{2}V$ and $\overline{\mathrm{v}^{\circ}(q)}\in\mathcal{Q}$ for $\mathrm{v}(s)\in S^{2}V/P^{\circ}$ and $\mathrm{v}^{\circ}(q)\in\mathcal{Q}/P$. If $(P_{t})$ is a smooth curve with $P_{0}=P$ and with derivative $\mathrm{v}$ at $t=0$, $s+t\overline{\mathrm{v}^{\circ}(p)}+o(t)\in P_{t}$ and $q+t\overline{\mathrm{v}(q)}+o(t)\in P^{\circ}_{t}$. Hence we get:
$\operatorname{Tr}\left(q\overline{\mathrm{v}^{\circ}(s)}+\overline{\mathrm{v}(q)}s\right)=0.$
An element $s\in P^{\circ}$ satisfies $s\in\operatorname{Ker}(\mathrm{v}^{\circ})$ if and only if $\operatorname{Tr}(\overline{\mathrm{v}(q)}s)=0$ for all $q\in\mathcal{Q}$, hence if and only if the corresponding linear form on $Q/P$ belongs to $\operatorname{Im}(\mathrm{v})^{\circ}$.
Now we prove the equivalence of the two definitions. If there exist $q]\in\operatorname{Im}(\mathrm{v})$ with $q$ positive, then for any $s\in\operatorname{Ker}(\mathrm{v}^{\circ})$, $\operatorname{Tr}(q\circ s)=0$. Hence $s$ is not a positive tensor. Therefore $\operatorname{Ker}(\mathrm{v}^{\circ})\subset P^{\circ}$ intersects trivially $S^{2}V^{\geq 0}$.
Conversely if $\operatorname{Ker}(\mathrm{v}^{\circ})$ intersects trivially $S^{2}V^{\geq 0}$, there exist $q\in\mathcal{Q}$ that do not vanish on $S^{2}V^{\geq 0}$, i.e. $q$ is positive. The class $q]\in\mathcal{Q}/P$ belongs to $\operatorname{Im}(\mathrm{v})$.